OpenStudy (anonymous):

what is the value of sin(-240)

OpenStudy (anonymous):

degree mode?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-240 + 360 = 120 so we need sin(120) What is sin (60)?

OpenStudy (anonymous):

Sin is positive in QII; thus, Sin (-240) = sin(120) = sin (60) = sqrt(3)/2

OpenStudy (anonymous):

so what do you use to solve these types of problems?

OpenStudy (anonymous):